A 1.3 kg block slides along a frictionless surface at 1.5 m/s . A second block, sliding at a faster 4.9 m/s , collides with the first from behind and sticks to it. The final velocity of the combined blocks is 2.0 m/s . What was the mass of the second block?
Answer #1

Using law of conservation of momentum,

(1.3 x 1.5) + (m x 4.9) = (1.3 + m) x 2

1.95 + 4.9 m = 2.6 + 2m

Mass, m = 0.224 kg

Comment in case any doubt please rate my answer..